OpenStudy (anonymous):

Which set of numbers does not create a triangle? 80 cm, 25 cm, 1 m 5 m, 12 m, 8 m 110 ft, 12 yd, 150 ft 7 yd, 13 yd, 19 yd

OpenStudy (anonymous):

ok i think this is how you do it. get all the units of the same so for the first one, change the 1m to cm. then plug it into the pathagorean therom.( a^2+b^2=c^2

OpenStudy (anonymous):

No one said that the triangle is right.

OpenStudy (anonymous):

ok nvm sorry!

OpenStudy (anonymous):

a, b, c are the three sides of a triangle if each side is less than the sum of the other tow sides and bigger than their diference.

OpenStudy (anonymous):

but you have to have the same units to compare right?

OpenStudy (anonymous):

@eliassaab It doesn't make a triangle?? The answer I originally got was 7 yd, 13 yd, 19 yd for some reason and I know the question wants to know which one does NOT make a triangle, just trying to understand :)

OpenStudy (anonymous):

ok well do what you did to find if it makes a triangle :). the one that doesnt work is not a triangle.

OpenStudy (anonymous):

@sparki1 I'm not sure how to tell if it makes a triangle tho :/ I haven't been able to learn this stuff yet sadly and it just was thrown at me today :/

OpenStudy (anonymous):

ok. ill try to figure it out for you.

OpenStudy (anonymous):

@sparki1 Thank you so much :)

OpenStudy (anonymous):

It is a series of three inequalities: a+b>c, a+c>b, and b+c>a use this. i think this is how you do it

OpenStudy (anonymous):

@sparki1 Omg thank you soooo much!! I get it now :) @eliassaab had the right answer! Thank you both so much I understand now :)

OpenStudy (anonymous):

7 yd, 13 yd, 19 yd can make a triangle 7 > 19-13=6